OpenStudy (ksaimouli):

√2tanxcosx−tanx=0 find all solutions (0,2pie)

OpenStudy (mr.math):

Didn't I just do this one?!

OpenStudy (ksaimouli):

\[tanxsin^2x=tanx\]

OpenStudy (ksaimouli):

plz give me answer i just want to verify

OpenStudy (mr.math):

Okay, we've got to the point \(\tan{x}=0 \text{ or }\cos{x}=\frac{1}{\sqrt{2}}\).

OpenStudy (ksaimouli):

ya i got but the radians

OpenStudy (ksaimouli):

0,2pie

OpenStudy (mr.math):

\(\tan{x}=0 \text{ at } x=0 \text{ and } x=\pi\).

OpenStudy (ksaimouli):

inbetween the circle

OpenStudy (mr.math):

\(\cos{x}=\frac{1}{\sqrt{2}} \text{ at } x=\frac{\pi}{4} \text{ or } \frac{3\pi}{4}.\)

OpenStudy (mr.math):

This is the solution of your first problem.

OpenStudy (ksaimouli):

coud u explain me second problem pz